I am starting a new job. They said they will pay me $650 on the 15th and the 30th of every month.I would make $1300 a month. There are different number of days in each month. How do I know what I make daily of each month?
Answer by nerdybill(7384)

.
Since you know you make $1300 per month
and, there are 12 months in a year
12 * 1300 = $15,600 per year
.
Since there are 52 weeks in a year:
15600/52 = $300 per week
.
Each week, you work only 5 days:
300/5 = $60 per day